Fill in table using the given formulas. S=n(n + 1)(2n + 1) I need to know what goes next to the rest of the numbers (3,4 &5). I am stuck.
n s
1 +1
2 n
3
4
5

Answer by stanbon(75887)

Fill in table using the given formulas. S=n(n + 1)(2n + 1) I need to know what goes next to the rest of the numbers (3,4 &5). I am stuck.
n... s
1....1(2)(3) = 6
2....2(3)(5) = 30
3....3(4)(7) = 84
4....5(5)(9) = 225
5....5(6)(11)= 30*11=330
=============================
Cheers,
Stan H.